# General Seemingly Unrelated Local Projections

### v2

@@ 2.3. Priors on impulse response functions
The joint prior thus allows for dependence across horizons $h$. Such a joint prior can in general have many parameters because we need to decide how much the impulse responses are correlated across horizons a priori, how much we should shrink towards our prior mean, whether this shrinkage should be horizon-specific, and so on. We now show how to set up a flexible joint prior that only depends on a few hyperparameters that need to be chosen.
@@
with prior means $\bm{\mu}_{\beta}=(\mu_{\beta 0},\ldots,\mu_{\beta\tilde{H}})^{\prime}$ and covariance matrix $\bm{V}_{\beta}$. These prior moments could, in principle, be directly chosen by the researcher as tuning parameters. For instance, setting $\bm{\mu}_{\beta}=\bm{0}$ and $\bm{V}_{\beta}$ such that its Cholesky factor is a lower triangular matrix with ones along and below its main diagonal implies that the prior on $\beta_{h}$ is centered on $\beta_{h-1}$. Alternatively, $\bm{\mu}_{\beta}$ can be set to have a particular shape or centered on the impulse responses of an auxiliary model such as a VAR (see Ferreira et al., 2023, for a similar approach) or using an empirical Bayes approach centered on the bias-adjusted IRF of Herbst and Johannsen (2024). In this case, $\bm{V}_{\beta}$ controls the weight put on $\bm{\mu}_{\beta}$ and becomes an important tuning parameter that needs to be chosen manually or by cross-validation.
An alternative route, which we adopt, is to treat both $\bm{\mu}_{\beta}$ and $\bm{V}_{\beta}$ as unknown and to use hierarchical priors to estimate them. This allows us to remain flexible, yet parsimonious, as we typically only need to choose a small set of prior hyperparameters that parameterize $\bm{\mu}_{\beta}$ and $\bm{V}_{\beta}$. We introduce a shrinkage prior by assuming that the prior covariance matrix is diagonal, $\bm{V}_{\beta}=\text{diag}(v_{\beta 0},\ldots,v_{\beta\tilde{H}})$. This allows us to test restrictions of the form $\beta_{h}=\mu_{\beta h}$ by setting the corresponding element on the diagonal of $\bm{V}_{\beta}$, $v_{\beta h}$, appropriately.
@@
We assume that $\bm{\mu}_{\beta}$ is a smooth function (in a sense we make precise below) and then let the prior variance $\bm{V}_{\beta}$ determine the weight to put on $\bm{\mu}_{\beta}$. Moreover, given that IRFs might feature certain characteristics (such as being hump-shaped or mean-reverting) we assume that $\bm{\mu}_{\beta}$ is a function of the horizons. Specifically, to impose smoothness on our prior impulse responses, we first define a function $\bm{\mu}_{\beta}(\bm{h})$ that takes a vector $\bm{h}=(0,\dots,\tilde{H})^{\prime}$ with real, non-negative entries as input. Although this might seem cumbersome at first since we will only use values $h=0,1,...,\tilde{H}$, this assumption allows us to borrow tools from the literature on Gaussian processes (GPs) which are popular in machine learning (for a textbook introduction, see Williams and Rasmussen, 2006). The function $\bm{\mu}_{\beta}(\bm{h})$ is modeled using GPs:
@@ 2.4. Illustration of the priors
To provide more intuition on how our GP prior works in practice, Figure 1 gives a few examples for different values of $\varsigma$ and $\xi$. The blue circles represent discrete observations of a dynamic multiplier $\bm{\beta}$, which we use to fit the GPs. Panels (a) to (d) always show the 95% credible intervals (gray shaded areas) of the prior (upper plot by panel) and the posterior (bottom plot by panel). The grey solid lines refer to three random samples from the prior, whereas for the posterior charts, we indicate the posterior median with a solid black line.
@@
Figure 1: Gaussian process prior and posterior distribution for various choices of $\varsigma$ and $\xi$ in panels (a) to (d). The posterior is fitted to a dynamic multiplier (blue circles). The gray shaded areas mark the 95% quantiles, darker gray lines are random draws from the prior while the solid black line indicates the posterior median.
@@
Figure 2: Variants of the global-local prior and implications for shrinkage. The top panels show the log of the marginal priors and the corresponding densities $p(\rho)$ of the shrinkage coefficients $\rho=1/(1+v_{\beta})$. The lower panels show samples and a few contour lines from the bivariate prior $p(\beta_{1}-\mu_{\beta 1},\beta_{2}-\mu_{\beta 2})$. Prior variants: NG with $\theta_{\lambda}=0.1$, $\tilde{\tau}^{2}\sim\mathcal{G}(0.1,0.1)$ in solid black; NG (fixed) with $\vartheta_{\lambda}=0.1$, $\tilde{\tau}^{2}=2$ in dashed grey; LASSO with $\theta_{\lambda}=1$, $\tilde{\tau}^{2}=2$ in dot-dashed grey; $\mathcal{N}$(0,1) indicates a standard normal prior in dotted grey.
@@
Next, we consider the shrinkage coefficient $\rho_{h}=1/(1+v_{\beta h})$. For illustrative purposes, we consider the simplified case $h=\tilde{H}=0$, and assume that $x_{t},u_{t}^{(0)}\sim\mathcal{N}(0,1)$ without additional control variables. [^15] The general expression is complicated by the presence of covariances across horizons and the variances of the shocks. We provide a discussion of this case in Appendix A; see also Polson and Scott (2010). The posterior mean $\overline{\beta}_{0}$ under these assumptions can be written as:
$\overline{\beta}_{0}=\rho_{0}\mu_{\beta 0}+(1-\rho_{0})\hat{\beta}_{0},$
that is, as a weighted average of the information represented in the data (where $\hat{\beta}_{0}$ is the least squares estimate) and the prior mean arising from the GP specification. If $v_{\beta 0}\to 0$, then $\rho_{0}\to 1$ and we thus end up with $\overline{\beta}_{0}=\mu_{\beta 0}$. By contrast, if $v_{\beta 0}\to\infty$ so that $\rho_{0}\to 0$ we end up with setting $\overline{\beta}_{0}=\hat{\beta}_{0}$.
With this in mind, consider the different shrinkage profiles induced by the various priors (except for the standard normal prior, which has a fixed variance and is thus excluded from this chart). The profile of the NG prior, which resembles the density of a Beta$(1/c,1/c)$ distribution (for $c$ being a large number), suggests cases that are either characterized by strong shrinkage so that $\beta_{h}\approx\mu_{\beta h}$ (corresponding to the case $\rho_{0}\approx 1$) or little shrinkage so that $\beta_{h}$ might deviate strongly from $\mu_{\beta h}$ (so that $\rho_{0}\approx 0)$. When we consider the LASSO the pole on 00 vanishes and we end up with cases in between. This indicates that the LASSO tends to overshrink significant effects (or in our framework to force $\beta_{h}$ towards $\mu_{\beta h}$ even if not supported by the data) whereas to induce too little shrinkage on cases where the data is consistent with the GP-induced prior. When we fix the $\tilde{\tau}^{2}$ we end up with a model that either implies very little shrinkage with a mode around $\rho_{0}=0.1$ or a lot of shrinkage (with a pole at $\rho_{0}=1$).
@@ 2.5. Priors on other model parameters
How we set the prior moments, $\bm{\mu}_{\gamma}=\text{vec}(\bm{M}_{\gamma})$ where $\bm{M}_{\gamma}$ is of size $k\times H$, and $\bm{V}_{\gamma}$, is inspired by the Minnesota-tradition. Specifically, we set the prior means to $1$ for the first own lag of $w_{t}$ (if the variable is in levels, in case it is in differences we set it to 00), and all remaining elements are zeroes. The informativeness with respect to the controls is set such that we have distinct hyperparameters associated with own and other lags as well as any deterministic variables; in addition, the prior tightness increases with the lag order.
Table 1 provides an overview of key tuning and hyperparameters. The joint posterior distribution of our framework is not available in closed form, which is why we use Markov chain Monte Carlo (MCMC) methods to sample from it. Most conditional distributions take a well-known form and are thus amenable to Gibbs sampling. Posteriors of parameters that do not follow any easy-to-sample from distributions are updated using Metropolis-Hastings updates. We provide details on posteriors and our full sampling algorithm in Appendix A.
@@ 3.1. Instrumenting shocks
$\begin{pmatrix}m_{t}\\ \bm{y}_{t}\end{pmatrix}=\begin{pmatrix}\phi\\ \bm{\beta}\end{pmatrix}x_{t}+\begin{pmatrix}\nu_{t}\\ \bm{u}_{t}\end{pmatrix},\quad\begin{pmatrix}\nu_{t}\\ \bm{u}_{t}\end{pmatrix}\sim\mathcal{N}\left(\bm{0},\begin{bmatrix}\sigma_{\nu}^{2}&\bm{0}^{\prime}\\ \bm{0}&\bm{\Sigma}_{u}\end{bmatrix}\right),$
@@ 3.3. Multiple shocks, their instruments, and measurement errors
where $\bm{X}_{t}=(\bm{I}_{H}\otimes\bm{x}_{t}^{\prime})$ and the shock-specific responses are stored in an $n_{x}\times H$-matrix $\bm{B}=(\bm{\beta}_{1},\ldots,\bm{\beta}_{n_{x}})^{\prime}$. The GP priors for the impulse responses, discussed for the single-shock case in Section 2.3, may be defined independently on the $\bm{\beta}_{i}$’s for $i=1,\ldots,n_{x},$ i.e., the rows of $\bm{B}$. Notice that this general approach also nests the case of a single instrument per shock by setting $n_{m}=1$. Moreover, the case that shock $i$ is observed is obtained by setting $n_{m}=1$, $\bm{\delta}_{i}=0$, $\bm{\phi}_{i}=1$ and Var$(\nu_{it})=0$. This can be achieved by choosing priors accordingly.
Another issue that commonly arises in applied work is that instruments for different shocks are correlated (see, e.g., Bruns et al., 2024). To control for this, Equation (18) can be modified as follows. Assuming, for simplicity, that $n_{m}=1$. Then, we obtain $\bm{\nu}_{t}=(\nu_{1t},\dots,\nu_{n_{x}t})^{\prime}$ which follows a Gaussian with zero mean but a (potentially) full covariance matrix $\bm{\Sigma}_{\nu}$. In the general case, we assume an informative inverse Wishart prior on $\bm{\Sigma}_{\nu}\sim\mathcal{W}^{-1}(s_{0\nu},\bm{S}_{0\nu})$.
@@
Indeed, in our empirical work we also use a special case of Equation (18) to extract a common component across shocks. This common component represents an interpretable shock (such as a common monetary shock present in multiple monetary policy instruments or a coordinated monetary/fiscal shock when studying monetary and fiscal policy jointly, as we do below). In that case, we propose one of two alternatives, reminiscent of how researchers have incorporated factors in Bayesian time series models: We either explicitly estimate the common factor from $n_{m}>1$ instruments by setting $n_{x}=1$, or we ex-ante compute the first principal component of all instruments and include that principal component as a joint instrument in our framework. Below we show results for both approaches. An extension to multiple common shocks is technically also feasible, but would likely lead to weak identification as multiple common shocks will likely be hard to disentangle. [^17] In general the data could be at least somewhat informative about whether or not the common component across instruments is due to noise or a meaningful structural shock as the inclusion of the common will change the fit of the LP equation.
@@ 3.4. Missing data
So far we have assumed that the dependent variable is observed for all relevant periods and horizons. This almost always leads a researcher to drop some data. We can use the LP structure to efficiently sample unobserved missing values of the dependent variable.
@@ 4. Step 4:
    • We may sample the missing values in $\bm{y}$ using the Gaussian distribution in Equation (19). Since most of the involved matrices are banded and/or sparse, we may update these quantities efficiently using precision sampling.

### v3

@@ 2.4. Priors on impulse response functions
The joint prior thus allows for dependence across horizons $h$. Such a joint prior can generally have many parameters because we need to decide how much the impulse responses are correlated across horizons a priori, how much we should shrink towards our prior mean, whether this shrinkage should be horizon-specific, and so on. We now show how to set up a flexible joint prior that only depends on a few hyperparameters that need to be chosen.
@@
[…]
@@
[…]
The function $\bm{\mu}_{\beta}(\bm{h})$ is modeled using GPs:
@@
Using an infinite-dimensional prior on $\bm{\mu}_{\beta}$ mitigates potential adverse effects arising from misspecification. Rather than imposing a rigid parametric structure, this approach permits the data to flexibly correct approximation errors introduced by the auxiliary likelihood. If misspecification is minor, the prior naturally shrinks $\bm{\mu}_{\beta}$ toward simpler functions. Conversely, it allows for greater complexity and adaptivity when deviations of the auxiliary likelihood from the true likelihood are substantial.
@@ 2.5. Illustration of the priors
To provide more intuition on how our GP prior works in practice, Figure 1 gives a few examples for different values of $\varsigma$ and $\xi$. The blue circles represent discrete observations of a dynamic multiplier $\bm{\beta}$, which we use to fit the GPs. Panels (a) to (d) always show the 95% credible intervals (gray shaded areas) of the prior (upper plot by panel) and the posterior (bottom plot by panel). The gray solid lines refer to three random samples from the prior, whereas for the posterior charts, we indicate the posterior median with a solid black line.
@@
Fig. 1: Gaussian process prior and posterior distribution for various choices of $\varsigma$ and $\xi$ in panels (a) to (d). The posterior is fitted to a dynamic multiplier (blue circles). The gray shaded areas mark the 95% quantiles, darker gray lines are random draws from the prior while the solid black line indicates the posterior median.
@@
Fig. 2: Variants of the global-local prior and implications for shrinkage. The top panels show the log of the marginal priors and the corresponding densities $p(\rho)$ of the shrinkage coefficients $\rho=1/(1+v_{\beta})$. The lower panels show samples and a few contour lines from the bivariate prior $p(\beta_{1}-\mu_{\beta 1},\beta_{2}-\mu_{\beta 2})$. Prior variants: NG with $\theta_{\lambda}=0.1$, $\tilde{\tau}^{2}\sim\mathcal{G}(0.1,0.1)$ in solid black; NG (fixed) with $\vartheta_{\lambda}=0.1$, $\tilde{\tau}^{2}=2$ in dashed gray; LASSO with $\theta_{\lambda}=1$, $\tilde{\tau}^{2}=2$ in dot-dashed gray; $\mathcal{N}$(0,1) indicates a standard normal prior in dotted gray.
@@
Next, we consider the shrinkage coefficient $\rho_{h}=1/(1+v_{\beta h})$. For illustrative purposes, we consider the simplified case $h=\tilde{H}=0$, and assume that $x_{t},u_{t}^{(0)}\sim\mathcal{N}(0,1)$ without additional control variables. [^14] The general expression is complicated by the presence of covariances across horizons and the variances of the shocks. We provide a discussion of this case in Appendix A; see also Polson and Scott (2010). The posterior mean $\overline{\beta}_{0}$ under these assumptions can be written as $\overline{\beta}_{0}=\rho_{0}\mu_{\beta 0}+(1-\rho_{0})\hat{\beta}_{0}$, that is, as a weighted average of the information represented in the data (where $\hat{\beta}_{0}$ is the least squares estimate) and the prior mean arising from the GP specification. If $v_{\beta 0}\to 0$, then $\rho_{0}\to 1$ and we thus end up with $\overline{\beta}_{0}=\mu_{\beta 0}$. By contrast, if $v_{\beta 0}\to\infty$ so that $\rho_{0}\to 0$ we end up with setting $\overline{\beta}_{0}=\hat{\beta}_{0}$.
With this in mind, consider the different shrinkage profiles induced by the various priors (except for the standard normal prior, which has a fixed variance and is thus excluded from this chart). The profile of the NG prior, which resembles the density of a Beta(1/c,1/c)(1/c,1/c) distribution (for $c$ being a large number), suggests cases that are either characterized by strong shrinkage so that $\beta_{h}\approx\mu_{\beta h}$ (corresponding to the case $\rho_{0}\approx 1$) or little shrinkage so that $\beta_{h}$ might deviate strongly from $\mu_{\beta h}$ (so that $\rho_{0}\approx 0)$. When we consider the LASSO the pole on 0 vanishes and we end up with cases in between. This indicates that the LASSO tends to overshrink significant effects (or in our framework to force $\beta_{h}$ towards $\mu_{\beta h}$ even if not supported by the data) whereas to induce too little shrinkage on cases where the data is consistent with the GP-induced prior. When we fix the $\tilde{\tau}^{2}$ we end up with a model that either implies very little shrinkage with a mode around $\rho_{0}=0.1$ or a lot of shrinkage (with a pole at $\rho_{0}=1$).
@@ 2.6. Priors on other model parameters
How we set the prior moments, $\bm{\mu}_{\gamma}=\text{vec}(\bm{M}_{\gamma})$ where $\bm{M}_{\gamma}$ is of size $k\times H$, and $\bm{V}_{\gamma}$, is inspired by the Minnesota-tradition. Specifically, we set the prior means to $1$ for the first own lag of $w_{t}$ (if the variable is in levels, in case it is in differences we set it to 0), and all remaining elements are zeroes. The informativeness with respect to the controls is set such that we have distinct hyperparameters associated with own and other lags as well as any deterministic variables; in addition, the prior tightness increases with the lag order.
@@
Table 1 provides an overview of key tuning and hyperparameters. The joint posterior distribution of our framework is not available in closed form, which is why we use Markov chain Monte Carlo (MCMC) methods to sample from it. Most conditional distributions take a well-known form and are thus amenable to Gibbs sampling. Posteriors of parameters that do not follow any easy-to-sample from distributions are updated using Metropolis-Hastings updates. We provide details on posteriors and our full sampling algorithm in Section 3.6 and Appendix A.
@@ 3.1. Instrumenting shocks
[…]
@@ 3.3. Multiple shocks, their instruments, and measurement errors
[…]
Another issue that commonly arises in applied work is that instruments for different shocks are correlated (see, e.g., Bruns et al., 2025). To control for this, Equation (18) can be modified as follows. Assuming, for simplicity, that $n_{m}=1$. Then, we obtain $\bm{\nu}_{t}=(\nu_{1t},\dots,\nu_{n_{x}t})^{\prime}$ which follows a Gaussian with zero mean but a (potentially) full covariance matrix $\bm{\Sigma}_{\nu}$. In the general case, we assume an informative inverse Wishart prior on $\bm{\Sigma}_{\nu}\sim\mathcal{W}^{-1}(s_{0\nu},\bm{S}_{0\nu})$.
@@
Indeed, in our empirical work we also use a special case of Equation (18) to extract a common component across shocks. This common component represents an interpretable shock (such as a common monetary shock present in multiple monetary policy instruments or a coordinated monetary/fiscal shock when studying monetary and fiscal policy jointly, as we do below). In that case, we propose one of two alternatives, reminiscent of how researchers have incorporated factors in Bayesian time series models: We either explicitly estimate the common factor from $n_{m}>1$ instruments by setting $n_{x}=1$, or we ex-ante compute the first principal component of all instruments and include that principal component as a joint instrument in our framework. Below we show results for both approaches. An extension to multiple common shocks is technically also feasible, but would likely lead to weak identification as multiple common shocks will be hard to disentangle. [^16] In general the data could be at least somewhat informative about whether or not the common component across instruments is due to noise or a meaningful structural shock as the inclusion of the common will change the fit of the LP equation.
@@ 3.4. Missing data
So far we have assumed that the dependent variable is observed for all relevant periods and horizons. However, with a fixed sample, longer horizon LPs have to be estimated using fewer observations to account for the lead of the left-hand side variable. We can use the LP structure to efficiently sample unobserved missing values of the dependent variable, circumventing this issue.
@@ 4. Step 4:
    • We may sample the missing values in $\bm{y}$ using the Gaussian distribution in Equation (19). Since most of the involved matrices are banded and/or sparse, we update these quantities efficiently using precision sampling.